OpenStudy (anonymous):

If sin Θ = 5 over 6, what are the values of cos Θ and tan Θ? cos Θ = ±11 over 36; tan Θ = ±1 over 11 cos Θ = ±square root 11 over 6; tan Θ = negative 1 over 11 cos Θ = ±11 over 6; tan Θ = negative 5 over 11 cos Θ = ±square root 11 over 6; tan Θ = ±5 square root 11 over 11

OpenStudy (johnweldon1993):

\[\large \sin(x) = \frac{opposite}{hypotenuse} = \frac{5}{6}\]

OpenStudy (johnweldon1993):

\[\large \cos(x) = \frac{adjacent}{hypotenuse}\] so you know whatever cos is...it will be a fraction with 6 in the denominator

OpenStudy (anonymous):

It is the last choice. I am taking this quiz too.

OpenStudy (anonymous):

really murphy23

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

yep but this is the only answer i know on the quiz lol

OpenStudy (anonymous):

can you see if you know any of the answers to my most recent questions?